Circular No. 5688

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505

MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

COMET MUELLER (1993a)

The following positions have been reported:

1993 UT α_{2000} δ_{2000} m_{1} Observer
Jan. 5.49317 9^{h}43^{m}05.^{s}65 +47^{o}51′57′′.4 15  Kojima
5.52089 9 43 03.80 +47 52 23.0  _{”}
5.52830 9 43 03.03 +47 52 30.0  _{”}
5.73438 9 42 47.47 +47 55 49.6  Sugie
5.74965 9 42 46.33 +47 56 05.6  _{”}
5.76615 9 42 45.05 +47 56 17.9  _{”}
5.77222 9 42 44.64 +47 56 26.8 14.5 Urata
5.77708 9 42 44.27 +47 56 30.1  _{”}
5.94084 9 42 31.96 +47 59 04.4  Pravec
5.95243 9 42 31.06 +47 59 17.2  _{”}
5.95493 9 42 31.07 +47 59 23.6  _{”}
5.97505 9 42 29.69 +47 59 41.3  _{”}
5.98145 9 42 28.68 +47 59 48.8  _{”}
5.98392 9 42 28.72 +47 59 51.4  _{”}
6.00059 9 42 27.50 +48 00 05.3  _{”}
6.00307 9 42 27.19 +48 00 07.2  _{”}
6.25706 9 42 07.92 +48 04 12.9  Balam
6.26480 9 42 07.18 +48 04 19.1  _{”}
6.27265 9 42 06.96 +48 04 27.7  _{”}
6.28007 9 42 06.00 +48 04 34.9  _{”}

T. Kojima (YGCO Chiyoda Observatory). 0.25-m f/3.4 Wright-Schmidt

camera. Communicated by S. Nakano. A. Sugie (Dynic Astronomical Observatory). 0.25-m f/3.4 Schmidt camera. T. Urata (Nihondaira Observatory Oohira station). 0.25-m f/3.4 hyper-

boloid astrocamera. P. Pravec (Ondrejov). 0.18-m f/5.6 Maksutov + CCD. D. D. Balam, J. B. Tatum, and R. M. Robb (Climenhaga Observatory,

Victoria). 0.5-m reflector + CCD. Measurer Balam. Comet outside of polygon of stars.

H. Mikuz, Ljubljana, Slovenia, reports that CCD photometry obtained with a 0.20-m f/2 Baker-Schmidt camera (+ V filter) shows the comet as having a 0′.6 coma of V = 14.0 and a highly condensed, almost starlike center, with a delicate 3′ fan-shaped tail in p.a. 180^{o}.

1993 January 7 Daniel W. E. Green